OpenStudy (anonymous):

Simplify: 8x^4y^0 /16xy a. x^3 /2y b. 1 /16xy c. 2x^3y d. x^5

OpenStudy (anonymous):

\[y^0=1\] so you have \[\frac{8x^4}{16xy}=\frac{x^3}{2y}\]

OpenStudy (anonymous):

Wait, what's the answer?

OpenStudy (anonymous):

Looks like a to me...